# Groups with few finite dimensional unitary representations

### v1

@@ Abstract.
The Bohr compactification of an arbitrary topological group $G$ is defined as the group compactification $(bG,b)$ with the following universal property: for every continuous homomorphism $h$ from $G$ into a compact group $K$ there is a continuous homomorphism $h^{b}$ from $bG$ into $K$ extending $h$ in the sense that $h=h^{b}\circ b$. The Bohr compactification $(bG,b)$ is the unique (up to equivalence) largest compactification of $G$. Although, for locally compact Abelian groups, the Bohr compactification is a big monster, for non-Abelian groups the situation is much more interesting and it can be said that all options are possible. Here we are interested in locally compact groups whose Bohr compactification is small. Among other results, we characterize when the Bohr the Bohr compactification of a locally compact group is topologically isomorphic to its Chu or unitary quasi-dual. Our results fixe some incorrect statements appeared in the literature.
@@ 1. Introduction
Suppose an algebraic group $G$ is equipped with two locally compact topologies $G_{1}=(G,\tau_{1})$ and $G_{2}=(G,\tau_{2})$. In case $G$ is abelian, we have that $G_{1}$ and $G_{2}$ are naturally (topologically) isomorphic if and only if so are their respective Bohr compactifications $bG_{1}$ and $bG_{2}$. Furthermore, if $\tau_{1}\subsetneq\tau_{2}$, then $|\frac{bG_{2}}{bG_{1}}|\geq 2^{\mathfrak{c}}$. Therefore the Bohr compactification of a locally compact abelian group completely characterizes its topological and algebraic structure.
This fact is known not to extend to non-Abelian groups and basically all options are possible for these groups. This happens because for a non-abelian group the set of finite-dimensional representations no longer reflects its overall topological and algebraic structure. In fact, there are compact groups that are the Bohr compactification of themselves when equipped with the discrete topology (the so-called self-bohrifying groups). Here we are interested in studying when the Bohr compactification of a locally compact group is close to itself. Our main result characterizes when the Bohr compactification of a locally compact group is topologically isomorphic to its Chu or unitary quasi-dual. We also provide an example that clarifies the application of the previous result. Our results fix some incorrect statements that have appeared in the literature.
@@ 2. Preliminaries
We also equip the subset $\textrm{irrep}_{n}(G)$, of all irreducible representations of dimension $n$, with the compact open topology inherited from $\textrm{rep}_{n}(G)$ and we denote by $\widehat{G}_{n}=\frac{\textrm{irrep}_{n}(G)}{\sim}=Q_{n}(\textrm{irrep}_{n}(G))$ the quotient space formed by the equivalence classes of irreducible representations of dimension $n$. Again, the space $\widehat{G}_{n}$ is equipped with the Fell topology defined above.
@@ 3. Self-bohrifying groups
If $G$ is a topological group, then $G_{d}$ denotes the same group with the discrete topology. According to Hart and Kunen [5], a compact group X is self-bohrifying if $(X,b)$ is the Bohr compactification of $X_{d}$, where $b\colon X_{d}\to X$ is the identity map. That is $b(X_{d})=X$. Previously, in [6, page 195], Comfort had defined a topological group $G$ to be a van der Waerden group ($vdW$-group) if every homomorphism from $G$ to a compact group is continuous. It is easily seen that a compact group is a $vdW$-group if and only if its topology is the finest totally bounded group topology on $G$. Finally, a topological group $G$ is tall if for each positive integer $n$ there are only finitely many classes $[\rho]\in\widehat{G}_{n}$. It is remarkable that all the notions introduced above are related for compact groups (cf. [5, Th. 1.15]).
@@
$(1)\Longrightarrow(2)$. By definition $\overline{G_{d}}=G$ if and only if every homomorphism of finite rank on $G$ is continuous. Moreover, the latter is equivalent to that every abstract subgroup of finite index in $G$ is open. On the other hand, $G_{d}$ is tall for representations with finite range if and only if it has only finitely many abstract subgroups of index n for each integer n. Thus, by [10, Th. 2], if $G_{d}$ is tall for representations with finite range, every abstract subgroup of finite index is open, which yields $(2)$.
$(2)\Longrightarrow(3)$. If $(3)$ is false, then for some $n\in\mathbb{N}$, the set $(\widehat{G_{d}})^{f}_{n}$ is infinite. According to [10, Th. 2], this means that some abstract subgroup of finite index is not open, which means that there are group homomorphisms of finite range defined on $G$ that are not continuous. Thus $(2)$ is also false.
@@ 4. Bohr compactification versus Chu duality
In this section we study when the Bohr compactification $bG$ of a locally compact group $G$ coincides with the Chu quasi-dual of $G$.
The main feature of Chu (or unitary) duality is the construction of a bidual of $G$ from $\textrm{rep}(G)$. This bidual is formed by the so-called quasi-representations. If we define $\mathbb{U}=\sqcup_{n<\omega}\mathbb{U}(n)$ (topological sum), a quasi-representation of $\textrm{rep}(G)$ is a continuous mapping $p:\textrm{rep}(G)\longrightarrow\mathbb{U}$ which preserves the main operations between unitary representations: direct sums, tensor products, unitary equivalence and sends the elements of $\textrm{rep}_{n}(G)$ into $\mathbb{U}(n)$ for all $n\in\mathbb{N}$ (see [1] or [7] for details). The set of all quasi-representations of $G$ equipped with the compact-open topology on $\textrm{rep}(G)$ is a topological group with pointwise multiplication as the composition law, called the Chu quasi-dual group of $G$ and denoted by $G^{xx}$. Thus, a neighborhood  base of the identity of $G^{xx}$ consists of sets of the form $[K_{n},V]=\{p\in G^{xx}:p(K_{n})\subset V\}$, where $V$ is any neighborhood  of the identity in $\mathbb{U}(n)$ and $K_{n}$ is any compact subset of $\textrm{rep}_{n}(G)$, $n\in\mathbb{N}$. It is easily verified that the evaluation map $\mathcal{E}_{G}:G\longrightarrow G^{xx}$ is a group homomorphism which is a monomorphism if and only if $G$ is MAP. We say that $G$ satisfies Chu duality when the evaluation map $\mathcal{E}_{G}$ is an isomorphism of topological groups. In this terminology, it was shown in [1], that LCA groups and compact groups satisfy Chu duality (indeed Chu duality reduces to Pontryagin duality and to Tannaka duality, respectively, for such groups). Here, the group $\mathcal{E}_{G}(G)$ is always assumed to be equipped with the topology inherited from $G^{xx}$. Remark that $G^{xx}_{|\textrm{rep}_{n}(G)}\subseteq C(\textrm{rep}_{n}(G),\mathbb{U}(n))$ , where $C(X,\mathbb{U}(n))$ denotes the space of all continuous functions from a topological space $X$ into $\mathbb{U}(n)$.
Next results provides a characterization of locally compact groups whose Chu quasi-dual coincides with its Bohr compactification. It fixes a wrong assertion stated in [3, Theorem 4.4] (see also [6, Proposition 2.2]).
@@
Proof. Necessity: Let $\mathcal{E}_{n}:\textrm{rep}_{n}(G)\longrightarrow{\rm rep}_{n}(G^{xx})$ the canonical evaluation map defined by $\mathcal{E}_{n}(D)(p)=p(D)$ for all $D\in\textrm{rep}_{n}(G)$ and $p\in G^{xx}$. We first check that $\mathcal{E}_{n}$ is continuous and injective. For the continuity, taking into account that the spaces $\textrm{rep}_{n}(G)$ and ${\rm rep}_{n}(G^{xx})$ are equipped with the compact open topology on $G$ and $G^{xx}$, respectively, let $\{D_{j}\}$ be a net converging to $D$ in $\textrm{rep}_{n}(G)$. Given an arbitrary compact subset $K$ of $G^{xx}$, since $\textrm{rep}_{n}(G)$ is locally compact and $G^{xx}_{|\textrm{rep}_{n}(G)}\subseteq C(\textrm{rep}_{n}(G),\mathbb{U}(n))$, it follows that $K$ is an equicontinuous subset of $C(\textrm{rep}_{n}(G),\mathbb{U}(n))$. Furthermore, the net $\{D_{j}\}$ converges to $D$ in the compact open topology on $\textrm{rep}_{n}(G)$ and every element $p\in G^{xx}$ is continuous on $\textrm{rep}_{n}(G)$, which means that the net $\{p(D_{j})\}$ converges to $p(D)$ for all $p\in G^{xx}$. That is to say, the net $\{\mathcal{E}_{n}(D_{j})\}$ converges to $\mathcal{E}_{n}(D)$ pointwise on $G^{xx}$. Hence, since $K$ is an equicontinuous subset of $C(\textrm{rep}_{n}(G),\mathbb{U}(n))$, it follows that $\{\mathcal{E}_{n}(D_{j})\}$ converges to $\mathcal{E}_{n}(D)$ uniformly on $K$. This yields the continuity of $\mathcal{E}_{n}$.
[…]
where $Q_{n}$ and $Q_{n}^{xx}$ are the canonical quotient mappings.
Therefore, let $K$ be an arbitrary compact subset of $\textrm{rep}_{n}(G)$. The continuity of $Q_{n}$ implies that $Q_{n}(K)$ will be a compact in $G^{x}_{n}$ and, by hypothesis, necessarily finite. Form a finite subset $\Phi_{n}=\{\sigma_{1},\ldots,\sigma_{k}\}\subseteq\textrm{rep}_{n}(G)$, taking a representative of each element in $K$. The unitary groups $\mathbb{U}(n)$ contain a neighborhood basis of the identity $I_{n}$ whose elements are invariant under conjugation. Furthermore, the density of $G$ in $G^{xx}$ implies that the elements in the latter group preserve conjugation. Altogether, this yields that the convergence on the finite set $\Phi_{n}$ is the same as uniform convergence on $K$, which completes the proof.
Next follows an example of a Chu reflexive discrete group $G$ whose partial dual spaces $\widehat{G}_{n}$, equipped with the Chu topology, are discrete.
Example 4.2.
Let $H$ be a finite simple non abelian group, let $I$ be an infinite index set and consider the discrete group $G=H^{(I)}$. Remark that the group $G$ also accepts a precompact topology as a dense subgroup of the compact group $H^{I}$. We want to prove that $(\widehat{G}_{n},t_{p}(G))$ is a discrete topological space. To see this, we show that for every $\varphi\in\textrm{irrep}_{n}(G)$ there is a finite subset $F$ of $G$ and $\epsilon>0$ such that if $\psi\in\textrm{irrep}_{n}(G)$ and $\psi\in N(\varphi,F,\epsilon)$, then $\psi\sim\varphi$.
@@
Suppose first that $\varphi(x)=(\varphi_{i_{1}}\circ\pi_{i_{1}})(x)$, $x=(x_{i})_{i\in I}\in G$, where $\pi_{i_{1}}\colon G\to H_{i_{1}}$ is the canonical projection of $G$ onto $H_{i_{1}}=H$. With some notational abuse, we identify $\varphi|_{F_{i_{1}}}$ with $\varphi_{i_{1}}$. Therefore we have $\varphi|_{F_{i_{1}}}=\varphi_{i_{1}}\in\textrm{irrep}_{n}(H)$. Take $F=F_{i_{1}}$, which is a finite subgroup of $G$, and $\epsilon=\frac{1}{n}$. It is known that if $\psi\in\textrm{irrep}_{n}(G)$ and $\|\varphi(x)-\psi(x)\|<\frac{1}{n}$ for all $x\in F=F_{i_{1}}\cong H$, then $\varphi_{i_{1}}=\varphi|_{F_{i_{1}}}\sim\psi|_{F_{i_{1}}}$ (see [7, Satz 12.1.6]). Moreover, assuming that $\psi\sim\psi_{j_{1}}\otimes\cdots\otimes\psi_{j_{M}}$, $\psi_{i_{l}}\in\textrm{irrep}_{m_{l}}(H)$, $1\leq l\leq M$ and $m_{1}\cdots m_{M}=n$. If $i_{1}\notin\{j_{1},\cdots,j_{M}\}$ then $\psi|_{F_{i_{1}}}\sim I_{n}$, where $I_{n}$ is the $n$-dimensional identity matrix. Therefore $\varphi_{i_{1}}\sim I_{n}$, which is impossible. Thus, there is an index $j_{k}=i_{1}$. Hence $\varphi|_{F_{i_{1}}}\sim\psi|_{F_{i_{1}}}$ yields
@@
By our induction hypothesis, and using [7, Satz 12.1.6], we deduce that
@@
It was proved in [1] that the discrete group $G=H^{(I)}$ is Chu reflexive. Nevertheless, we include a short version of it here for the reader’s sake. Let $\varphi\in\textrm{rep}_{n}(H)$ be a faithful unitary representation of $H$ and set $\varphi_{i}=\varphi\circ\pi_{i}$, where $\pi_{i}\colon G\to H_{i}$ is the canonical projection of $G$ onto $H_{i}=H$. Since $G=H^{(I)}$ is a discrete group, it is easyly verified that the trivial representation $I_{n}$ (the one that sends every element to $I_{n}$) is a closure point of the discrete subset $\{\varphi_{i}:i\in J\}$ for every infinite subset $J\subseteq I$. On the other hand, we have that $bG\cong H^{I}$ (see [1, Cor. 3.10]). Thus, ir order to show that $G$ is Chu reflexive, it will suffice to no element in $H^{I}\setminus G$ defines a continuous quasi-representation. Hence, take $(x_{i})\in H^{I}\setminus G$. Then there is an infinite subset $J\subseteq I$ such that $x_{i}\not=e_{H}$ for all $i\in J$. Furthermore, since the group $H$ is finite, we may assume without loss of generality that $x_{i}=x_{j}$ for all $i,j$ in $J$. As a consequence, there is $A\in\mathbb{U}_{n}$ such that $(x_{i})[\varphi_{i}]=\varphi(x_{i})=A\not=I_{n}$ for all $i\in J$. This implies that $(x_{i})$ is not continuous on $\textrm{rep}_{n}(G)$, which completes the proof.
@@
It was stated in [3, Theorem 4.4] that if $G$ is a locally compact group such that $\widehat{G}_{n}$ is discrete for all $n\in\mathbb{N}$, then $bG=G^{xx}$ and therefore the group $G$ may not be Chu reflexive. This result is incorrect as we have shown in Example 4.2 and the reason this happens is subtle. It was proven in [2, Cor. 5.6] that if $G$ is a discrete group and some partial dual object $\widehat{G}_{n}$ is infinite, then the trivial representation $1_{G}$ is not an isolated point in in some $\widehat{G}_{m}\cup\{1_{G}\}$ with $1\leq m\leq n^{2}$. This means that although the topological space $\widehat{G}_{m}$ may be discrete, the uniformity that $\widehat{G}_{m}$ inherits from $\widehat{G}$ is not, as asserted in [3]. This said, we want to emphasize that the main consequences of [3, Theorem 4.4] are preserved without modification using Theorem 4.1. In particular, we have:

### v2

# Bohr compactification and Chu duality of non-abelian locally compact groups

@@ Abstract.
The Bohr compactification of an arbitrary topological group $G$ is defined as the group compactification $(bG,b)$ with the following universal property: for every continuous homomorphism $h$ from $G$ into a compact group $K$ there is a continuous homomorphism $h^{b}$ from $bG$ into $K$ that extends $h$ in the sense that $h=h^{b}\circ b$. The Bohr compactification $(bG,b)$ is the unique (up to equivalence) largest compactification of $G$. Although, for locally compact abelian groups the Bohr compactification is a big monster, for non-Abelian groups the situation is much more interesting and one can say that all options are possible. Here we discuss the relation between the notions of duality and the Bohr compactification of non-Abelian locally compact groups. Our main result establishes the existence of noncompact Chu reflexive groups whose partial dual spaces $\widehat{G}_{n}$, equipped with the Fell topology, are discrete. Furthermore, we characterize when the Bohr compactification of a locally compact group is topologically isomorphic to its Chu or unitary quasi-dual. These results fix some incorrect statements that have appeared in the literature.
@@ 1. Introduction
Suppose an algebraic group $G$ is equipped with two locally compact topologies $G_{1}=(G,\tau_{1})$ and $G_{2}=(G,\tau_{2})$. In case $G$ is abelian, we have that the identity mapping of ${\rm i_{d}}\colon G_{1}\to G_{2}$ is a topological group isomorphism if and only if there is a topological isomorphism $\Phi\colon bG_{1}\to bG_{2}$ such that $\Phi_{|G}={\rm i_{d}}$, where $bG_{1}$ and $bG_{2}$ denote the the Bohr compactifications of $G_{1}$ and $G_{2}$, respectively. Furthermore, if $\tau_{1}\subsetneq\tau_{2}$, then $|\frac{bG_{2}}{bG_{1}}|\geq 2^{\mathfrak{c}}$. Therefore the Bohr compactification of a locally compact abelian group completely characterizes its topological and algebraic structure.
This fact does not extend to non-Abelian groups, where essentially all options are possible. This occurs because, for a non-Abelian group, the set of finite-dimensional representations no longer reflects its overall topological and algebraic structure. In fact, there exist compact groups that are the Bohr compactification of themselves when equipped with the discrete topology (these are known as self-Bohrifying groups).
In this project, we aim to clarify the relationship between the Bohr compactification of locally compact groups and the Chu (or unitary) duality of these groups. For example, we characterize the conditions under which the Bohr compactification of a locally compact group is topologically isomorphic to its Chu or unitary quasi-dual.
We also address the question of whether a group whose Chu dual is discrete must necessarily be compact. Specifically, we establish the existence of non-compact Chu-reflexive groups whose partial dual spaces, $\widehat{G}_{n}$, equipped with the Fell topology, are discrete. Our results correct some inaccurate statements found in the literature.
@@ 2. Preliminaries
We also equip the subset $\textrm{irrep}_{n}(G)$, of all irreducible representations of dimension $n$, with the compact open topology inherited from $\textrm{rep}_{n}(G)$ and we denote by $\widehat{G}_{n}=\frac{\textrm{irrep}_{n}(G)}{\sim}=Q_{n}(\textrm{irrep}_{n}(G))$ the quotient space formed by the equivalence classes of irreducible representations of dimension $n$. Again, the space $\widehat{G}_{n}$ is equipped with the Fell topology defined above. Thus the partial dual object $\widehat{G}_{n}$ is a topological subspace of $G_{n}^{x}$ for all $n\in{\mathbb{N}}$.
@@ 3. Self-bohrifying groups
If $G$ is a topological group, then $G_{d}$ denotes the same group with the discrete topology. According to Hart and Kunen [6], a compact group X is self-bohrifying if $(X,b)$ is the Bohr compactification of $X_{d}$, where $b\colon X_{d}\to X$ is the identity map. That is $b(X_{d})=X$. Previously, in [6, page 195], Comfort had defined a topological group $G$ to be a van der Waerden group ($vdW$-group) if every homomorphism from $G$ to a compact group is continuous. It is easily seen that a compact group is a $vdW$-group if and only if its topology is the finest totally bounded group topology on $G$. Finally, a topological group $G$ is tall if for each positive integer $n$ there are only finitely many classes $[\rho]\in\widehat{G}_{n}$. It is remarkable that all the notions introduced above are related for compact groups (cf. [6, Th. 1.15]).
@@
$(1)\Longrightarrow(2)$. By definition $\overline{G_{d}}=G$ if and only if every homomorphism of finite rank on $G$ is continuous. Moreover, the latter is equivalent to that every abstract subgroup of finite index in $G$ is open. On the other hand, $G_{d}$ is tall for representations with finite range if and only if it has only finitely many abstract subgroups of index n for each integer n. Thus, by [11, Th. 2], if $G_{d}$ is tall for representations with finite range, every abstract subgroup of finite index is open, which yields $(2)$.
$(2)\Longrightarrow(3)$. If $(3)$ is false, then for some $n\in\mathbb{N}$, the set $(\widehat{G_{d}})^{f}_{n}$ is infinite. According to [11, Th. 2], this means that some abstract subgroup of finite index is not open, which means that there are group homomorphisms of finite range defined on $G$ that are not continuous. Thus $(2)$ is also false.
@@ 4. Bohr compactification versus Chu duality
In the previous secction we have seen how selfborifying groups are characterized by the property that their partial duals $\widehat{G}_{n}$ are finite for all $n\in{\mathbb{N}}$. In this section, we explore when the Bohr compactification of a locally compact group $G$ coincides with its Chu or unitary bidual. The main feature of Chu duality is the construction of a bidual of $G$ from $\textrm{rep}(G)$. This bidual is formed by the so-called quasi-representations. If we define $\mathbb{U}=\sqcup_{n<\omega}\mathbb{U}(n)$ (topological sum), a quasi-representation of $\textrm{rep}(G)$ is a continuous mapping $p:\textrm{rep}(G)\longrightarrow\mathbb{U}$ which preserves the main operations defined between unitary representations: direct sums, tensor products, unitary equivalence and sends the elements of $\textrm{rep}_{n}(G)$ into $\mathbb{U}(n)$ for all $n\in\mathbb{N}$ (see [1] or [8] for details). The set of all quasi-representations of $\textrm{rep}(G)$ equipped with the compact-open topology on $\textrm{rep}(G)$ is a topological group with pointwise multiplication as the composition law, called the Chu bidual group of $G$ and denoted by $G^{xx}$. Thus, a neighborhood  base of the identity of $G^{xx}$ consists of sets of the form $[K_{n},V]=\{p\in G^{xx}:p(K_{n})\subset V\}$, where $V$ is any neighborhood  of the identity in $\mathbb{U}(n)$ and $K_{n}$ is any compact subset of $\textrm{rep}_{n}(G)$, $n\in\mathbb{N}$. It is easily verified that the evaluation map $\mathcal{E}_{G}:G\longrightarrow G^{xx}$ is a group homomorphism which is a monomorphism if and only if $G$ is MAP. We say that $G$ satisfies Chu duality when the evaluation map $\mathcal{E}_{G}$ is an isomorphism of topological groups. In this terminology, it was shown in [1], that LCA groups and compact groups satisfy Chu duality (indeed Chu duality reduces to Pontryagin duality and to Tannaka duality, respectively, for such groups). Here, the group $\mathcal{E}_{G}(G)$ is always assumed to be equipped with the topology inherited from $G^{xx}$. Remark that $G^{xx}_{|\textrm{rep}_{n}(G)}\subseteq C(\textrm{rep}_{n}(G),\mathbb{U}(n))$ , where $C(X,\mathbb{U}(n))$ denotes the space of all continuous functions from a topological space $X$ into $\mathbb{U}(n)$.
Our first result demonstrates the existence of non compact Chu reflexive groups whose partial dual spaces $\widehat{G}_{n}$, equipped with the Fell topology, are discrete. Theorems 4.1 and 4.2 fix a wrong assertion stated in [4, Theorem 4.4] (see also [7, Proposition 2.2]).
@@
Let $H$ be a finite simple non abelian group, let $I$ be an infinite index set and consider the discrete group $G=H^{(I)}$. Then $G$ is Chu reflexive and $(\widehat{G}_{n},t_{p}(G))$ is a discrete topological space for all $n\in{\mathbb{N}}$.
Proof.
Remark that the group $G$ also accepts a precompact topology as a dense subgroup of the compact group $H^{I}$. We want to prove that $(\widehat{G}_{n},t_{p}(G))$ is a discrete topological space. To see this, we show that for every $\varphi\in\textrm{irrep}_{n}(G)$ there is a finite subset $F$ of $G$ and $\epsilon>0$ such that if $\psi\in\textrm{irrep}_{n}(G)$ and $\psi\in N(\varphi,F,\epsilon)$, then $\psi\sim\varphi$.
@@
Suppose first that $\varphi(x)=(\varphi_{i_{1}}\circ\pi_{i_{1}})(x)$, $x=(x_{i})_{i\in I}\in G$, where $\pi_{i_{1}}\colon G\to H_{i_{1}}$ is the canonical projection of $G$ onto $H_{i_{1}}=H$. With some notational abuse, we identify $\varphi|_{F_{i_{1}}}$ with $\varphi_{i_{1}}$. Therefore we have $\varphi|_{F_{i_{1}}}=\varphi_{i_{1}}\in\textrm{irrep}_{n}(H)$. Take $F=F_{i_{1}}$, which is a finite subgroup of $G$, and $\epsilon=\frac{1}{n}$. It is known that if $\psi\in\textrm{irrep}_{n}(G)$ and $\|\varphi(x)-\psi(x)\|<\frac{1}{n}$ for all $x\in F=F_{i_{1}}\cong H$, then $\varphi_{i_{1}}=\varphi|_{F_{i_{1}}}\sim\psi|_{F_{i_{1}}}$ (see [8, Satz 12.1.6]). Moreover, assuming that $\psi\sim\psi_{j_{1}}\otimes\cdots\otimes\psi_{j_{M}}$, $\psi_{i_{l}}\in\textrm{irrep}_{m_{l}}(H)$, $1\leq l\leq M$ and $m_{1}\cdots m_{M}=n$. If $i_{1}\notin\{j_{1},\cdots,j_{M}\}$ then $\psi|_{F_{i_{1}}}\sim I_{n}$, where $I_{n}$ is the $n$-dimensional identity matrix. Therefore $\varphi_{i_{1}}\sim I_{n}$, which is impossible. Thus, there is an index $j_{k}=i_{1}$. Hence $\varphi|_{F_{i_{1}}}\sim\psi|_{F_{i_{1}}}$ yields
@@
By our induction hypothesis, and using [8, Satz 12.1.6], we deduce that
@@
It was proved in [1] that the discrete group $G=H^{(I)}$ is Chu reflexive. Nevertheless, we include a short version of it here for the reader’s sake. Let $\varphi\in\textrm{rep}_{n}(H)$ be a faithful unitary representation of $H$ and set $\varphi_{i}=\varphi\circ\pi_{i}$, where $\pi_{i}\colon G\to H_{i}$ is the canonical projection of $G$ onto $H_{i}=H$. Since $G=H^{(I)}$ is a discrete group, it is easyly verified that the trivial representation $I_{n}$ (the one that sends every element to $I_{n}$) is a closure point of the discrete subset $\{\varphi_{i}:i\in J\}$ for every infinite subset $J\subseteq I$. On the other hand, we have that $bG\cong H^{I}$ (see [1, Cor. 3.10]). Thus, ir order to show that $G$ is Chu reflexive, it will suffice to verify that no element in $H^{I}\setminus G$ defines a continuous quasi-representation. Hence, take $(x_{i})\in H^{I}\setminus G$. Then there is an infinite subset $J\subseteq I$ such that $x_{i}\not=e_{H}$ for all $i\in J$. Furthermore, since the group $H$ is finite, we may assume without loss of generality that $x_{i}=x_{j}$ for all $i,j$ in $J$. As a consequence, there is $A\in\mathbb{U}_{n}$ such that $(x_{i})[\varphi_{i}]=\varphi(x_{i})=A\not=I_{n}$ for all $i\in J$. This implies that $(x_{i})$ is not continuous on $\textrm{rep}_{n}(G)$, which completes the proof. ∎
The following result completes Theorem 4.1 by giving a characterization of locally compact groups whose Chu quasi-dual coincides with their Bohr compactification.
Theorem 4.2.
Proof.
Necessity: Let $\mathcal{E}_{n}:\textrm{rep}_{n}(G)\longrightarrow{\rm rep}_{n}(G^{xx})$ the canonical evaluation map defined by $\mathcal{E}_{n}(D)(p)=p(D)$ for all $D\in\textrm{rep}_{n}(G)$ and $p\in G^{xx}$. We first check that $\mathcal{E}_{n}$ is continuous and injective. For the continuity, taking into account that the spaces $\textrm{rep}_{n}(G)$ and ${\rm rep}_{n}(G^{xx})$ are equipped with the compact open topology on $G$ and $G^{xx}$, respectively, let $\{D_{j}\}$ be a net converging to $D$ in $\textrm{rep}_{n}(G)$. Given an arbitrary compact subset $K$ of $G^{xx}$, since $\textrm{rep}_{n}(G)$ is locally compact and $G^{xx}_{|\textrm{rep}_{n}(G)}\subseteq C(\textrm{rep}_{n}(G),\mathbb{U}(n))$, it follows that $K$ is an equicontinuous subset of $C(\textrm{rep}_{n}(G),\mathbb{U}(n))$. Furthermore, the net $\{D_{j}\}$ converges to $D$ in the compact open topology on $\textrm{rep}_{n}(G)$ and every element $p\in G^{xx}$ is continuous on $\textrm{rep}_{n}(G)$, which means that the net $\{p(D_{j})\}$ converges to $p(D)$ for all $p\in G^{xx}$. That is to say, the net $\{\mathcal{E}_{n}(D_{j})\}$ converges to $\mathcal{E}_{n}(D)$ pointwise on $G^{xx}$. Hence, since $K$ is an equicontinuous subset of $C(\textrm{rep}_{n}(G),\mathbb{U}(n))$, it follows that $\{\mathcal{E}_{n}(D_{j})\}$ converges to $\mathcal{E}_{n}(D)$ uniformly on $K$. This yields the continuity of $\mathcal{E}_{n}$.
[…]
where $Q_{n}$ and $Q_{n}^{xx}$ are the canonical quotient mappings.
Therefore, let $K$ be an arbitrary compact subset of $\textrm{rep}_{n}(G)$. The continuity of $Q_{n}$ implies that $Q_{n}(K)$ will be a compact in $G^{x}_{n}$ and, by hypothesis, necessarily finite. Form a finite subset $\Phi_{n}=\{\sigma_{1},\ldots,\sigma_{k}\}\subseteq\textrm{rep}_{n}(G)$, taking a representative of each element in $K$. The unitary groups $\mathbb{U}(n)$ contain a neighborhood basis of the identity $I_{n}$ whose elements are invariant under conjugation. Furthermore, the density of $G$ in $G^{xx}$ implies that the elements in the latter group preserve conjugation. Altogether, this yields that the convergence on the finite set $\Phi_{n}$ is the same as uniform convergence on $K$, which completes the proof. ∎
The next result improves Theorem 4.2 for discrete groups, clarifies Theorem 4.1, and extends Proposition 3.1.
Corollary 4.3.
Let $G$ be a discrete group. Then $G^{xx}$ is topologically isomorphic to $bG$ if and only if the space $\widehat{G}_{n}$ is finite for any $n\in\mathbb{N}$.
Proof.
Necessity: Since $G$ is discrete, the dual object $\widehat{G}$ is quasicompact (see [2, §$\S$18.1]). On the other hand, the partial dual $\widehat{G}_{n}$ is a topological subspace of $G^{x}_{n}$ that is discrete by Theorem 4.2. As $\bigcup\limits_{i=1}^{n}\widehat{G}_{i}$ is closed in $\widehat{G}$, it is also quasicompact and discrete, hence finite.
Sufficiency: As in Theorem 4.2, in order to prove that $bG=G^{xx}$ it suffices to prove that $G^{xx}$ is precompact. Now, since $\widehat{G}_{m}$ is finite, it follows that the quotient uniformity canonically associated to $\widehat{G}_{m}$ by the quotient mapping $Q_{m}$ is the discrete uniformity. Hence, the subset $\Lambda_{n}:=\{\sigma_{i}\}_{i\in I_{n}}\subset\textrm{irrep}_{n}(G)$, formed by choosing a representative of each equivalence class, is finite and, hence, inherites the discrete uniformity from $\textrm{rep}_{n}(G)$. We next show how the topology of uniform convergence on compact subsets of $\textrm{rep}_{n}(G)$ and the topology of pointwise convergence on $\Lambda_{n}$ coalesce. Then the proof is completed by observing that the latter topology is trivially precompact.
Indeed, choose an arbitrary compact subset $K\subseteq\textrm{rep}_{n}(G)$. Each element of $K$ decomposes as the direct sum of a finite number of irreducible subrepresentations, let’s denote by $[K]_{m}\subseteq\textrm{irrep}_{m}(G)$ the subset of all irreducible components of elements of $K$ of dimension $m$. By our initial assumption, we have that $\mathbb{Q}_{m}([K]_{m})$ is a finite subset of $\widehat{G}_{m}$ for $1\leq m\leq n$. Form a finite subset $\{\sigma_{1},\ldots,\sigma_{k}\}$ of $\Lambda$ taking a representative of each element in $\Phi_{n}=\bigcup_{m=1}^{n}\Psi_{m}([K]_{m})$. Arguing as in the last paragraph of Theorem 4.2, we conclude that pointwise convergence on the finite set $\Phi_{n}$ is the same as uniform convergence on $K$. ∎
@@
It was stated in [4, Theorem 4.4] that if $G$ is a locally compact group such that $\widehat{G}_{n}$ is discrete for all $n\in\mathbb{N}$, then $bG=G^{xx}$ and therefore the group $G$ may not be Chu reflexive. This result is incorrect as we have shown in Theorem 4.1 and the reason why this happens is subtle. It was proven in [3, Cor. 5.6] that if $G$ is a discrete group and some partial dual object $\widehat{G}_{n}$ is infinite, then the trivial representation $1_{G}$ is not an isolated point in in some $\widehat{G}_{m}\cup\{1_{G}\}$ with $1\leq m\leq n^{2}$. This means that although the topological space $\widehat{G}_{m}$ may be discrete, the uniformity that $\widehat{G}_{m}$ inherits from $\widehat{G}$ is not, as asserted in [4]. This said, we want to emphasize that the main consequences of [4, Theorem 4.4] are preserved without modification using Theorem 4.2. In particular, we have: